The expected total value after playing 40 rounds of roulette betting $10 on one number each time is

Select one:

a. 0.0

b. -15.0

c. -21.05

d. +15.0

e. +21.2
Answer #1

here as probability of winning =18/38

therefore expected winning on one round =(18/38)*10-(20/38)*10=-0.526

hence for 40 rounds ; expected value =40*(-0.526)=-21.05

option C is correct